OpenStudy (anonymous):

. Suppose $8,900 is invested in an account at an annual interest rate of 5.3% compounded continuously. How long (to the nearest tenth of a year) will it take the investment to double in size?

jimthompson5910 (jim_thompson5910):

\[\Large A = Pe^{rt}\] \[\Large A = 8900e^{0.053t}\] \[\Large 2*8900 = 8900e^{0.053t}\] \[\Large 2 = e^{0.053t}\] \[\Large \ln(2) = 0.053t\] \[\Large t = \frac{\ln(2)}{0.053}\] \[\Large t = ???\]

OpenStudy (anonymous):

I got t=13.08? I have a really hard time with solving word problems.

jimthompson5910 (jim_thompson5910):

you got it, so the final answer rounds to 13.1 years

OpenStudy (anonymous):

Thanks so much for your help!

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

Do you happen to know the formula for this problem? . A candidate’s share of the votes varies directly as the number of votes cast for the candidate. In a city council election with four candidates, Al received 540 votes and earned a 36% share of the votes. Bob received 720 votes. What was Bob’s corresponding share of the votes?

jimthompson5910 (jim_thompson5910):

it sounds like a ratio problem 540/36 = 720/x solve that for x to get your answer